Given the demand curve for commodity as Q=60-3p where Q is the quantity demanded and P the price per unit . Determine the price elasticity of demand at P= 10


1
Expert's answer
2021-06-09T17:55:07-0400

Solution:

Price elasticity of demand = "\frac{\%\triangle Q}{\%\triangle P} = \frac{\triangle Q}{\triangle P} \times \frac{P}{Q}"


The slope of Q = a – bP


"\frac{\triangle Q}{\triangle P} = -b"


Price elasticity of demand = "-b \times \frac{P}{Q}"


Derive quantity by substituting the price value, when P = 10:

Q = 60 – 3P

Q = 60 – 3(10)

Q = 60 – 30 = 30

Q = 30

P = 10

-b = -3

Therefore:

Price elasticity of demand "= -3 \times \frac{10}{30} = -1"

Price elasticity of demand = 1